A propeller is modeled as five identical uniform rods extending radially from its axis. The length and mass of each rod are 0.845 m and 2.79 kg, respectively. When the propellor rotates at 511 rpm (revolutions per minute), what is its rotational kinetic energy?
Answer #1

m =2.79 kg, L = 0.845 m

\omega =511 rev/min = 511x(2\pi/60)

\omega = 53.48 rad/s

I = 5xML2/12

Rotatinal kinetic energy K.E =(1.2)\omega2

K.E =(1/2)(5xML2/12)\omega2

K.E =(1/2) [5x2.79x0.845x0.845/12](53.48)2

K.E = 1187 J
Answer #2

The propeller's rotational kinetic energy K is

K=12Iω2

where I is its moment of inertia with respect to the axis of rotation and ω is its angular speed.

The propeller consists of five identical uniform rods rotating about their ends. Its moment of inertia I is therefore five times the moment of inertia of a single uniform rod of mass m and length L with respect to a perpendicular axis passing through its end.

I=5(13mL2)=53mL2

The rotational kinetic energy K of the propeller is therefore

K=12(53mL2)ω2=56mL2ω2

in terms of given quantities, although ω is given in rpm rather than in the appropriate SI unit of rad/s. Convert ω.ω=(515 rev/min)(2π rad1 rev)(1 min60 s)=53.9 rad/s

Now, substitute the data in the expression for K in order to obtain the numerical answer.




answered by: Muhammad Aslam